Consider a cavity consisting of two plane mirrors separated by a distance 60 cm in air. 

Calculate the mode number corresponding to the wavelength λ=600 nm. Also 

calculate the frequency spacing between the two longitudinal modes.


1
Expert's answer
2021-01-13T11:42:01-0500

"\Delta V_m=\frac{C}{2L}=\frac{\left(3\times10^8ms^{-1}\right)}{0.6\times2}=2.5\times10^8Hz"

"mode\:number=\frac{\Delta V}{\Delta V_m}=\frac{5\times10^{14}Hz}{2.5\times10^8Hz}=2\times10^6mods"

"Frequency\:of\:spacing\left(\Delta V\right)=\frac{Velocity\:of\:air}{Wavelength}=\frac{3\times10^8ms^{-1}}{600\times10^{-9m}}=5\times10^{14}Hz"